# What Does It Take to Be a Good AI Research Agent? Studying the Role of Ideation Diversity

### v1

@@ A.1 Controlling Ideation Diversity Using Temperature
With the intuition that the temperature sampling parameter has an impact on ideation diversity and performance, we run several experiments intervening on the sampling temperature. To isolate the impact of temperature, we use $\text{AIRA}_{\text{Greedy}}$ but remove all mechanisms that enhance ideation diversity (sibling memory, prompt-adaptive complexity, mention of diversity in the system prompt), as described in section 3.3, and experiment with different temperatures, above and below the recommended temperature of 0.6 for DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability).
Figure 12 shows that changing temperature does not have an impact on performance (neither beneficial nor detrimental), assessed as medal rate. Results hold across alternative metrics like valid submission rate, average normalized score, and percentile, except for Elo, the only metric where increased temperature significantly leads to improved performance. Despite our effort in isolating ideation diversity, we hypothesize that the reason for these negative results is that modifying the temperature parameter (instead of using the recommended one) also affects the agent in additional manners (Renze_2024; wu2025roletemperaturesamplingtesttime; li2025exploringimpacttemperaturelarge). For example, we would expect the implementation capabilities to also be affected, [^2] Surprisingly, Renze_2024 found no statistically significant effect in problem solving skills of LLMs for certain temperature ranges. and there could be second-order effects that are hard to reason about. We leave further investigating the temperature-based results as future work.
@@ A.3 Models Used By Agents For Image Classification Tasks
Focusing on the neural architectures used by the agents in image classification tasks in Figure 14, [^3] Image classification is the largest category in MLE-bench (8 tasks out of 22 in MLE-bench lite) we observe that AIDE relies on the EfficientNet architecture (and its variants) for almost 40% of the tasks. For 75% of its attempts, AIDE uses only 3 different architectures: EfficientNet, ResNet (he2016deep) and LightGBM. $\text{AIRA}_{\text{Greedy}}$ uses a wider range of architectures, with EfficientNet, ConvNeXt (liu2022convnet), and ViT (dosovitskiy2020image) making only 38% of the agents’ initial ideas.

### v2

@@ A.1 Controlling Ideation Diversity Using Temperature
With the intuition that the temperature sampling parameter has an impact on ideation diversity and performance, we run several experiments intervening on the sampling temperature. To isolate the impact of temperature, we use $\text{AIRA}_{\text{Greedy}}$ but remove all mechanisms that enhance ideation diversity (sibling memory, prompt-adaptive complexity, mention of diversity in the system prompt), as described in section 3.3, and experiment with different temperatures, above and below the recommended temperature of 0.6 for DeepSeek-R1 (DeepSeek-AI et al., 2025).
Figure 12 shows that changing temperature does not have an impact on performance (neither beneficial nor detrimental), assessed as medal rate. Results hold across alternative metrics like valid submission rate, average normalized score, and percentile, except for Elo, the only metric where increased temperature significantly leads to improved performance. Despite our effort in isolating ideation diversity, we hypothesize that the reason for these negative results is that modifying the temperature parameter (instead of using the recommended one) also affects the agent in additional manners (Renze, 2024; Wu et al., 2025; Li et al., 2025b). For example, we would expect the implementation capabilities to also be affected, [^2] Surprisingly, Renze (2024) found no statistically significant effect in problem solving skills of LLMs for certain temperature ranges. and there could be second-order effects that are hard to reason about. We leave further investigating the temperature-based results as future work.
@@ A.3 Models Used By Agents For Image Classification Tasks
Focusing on the neural architectures used by the agents in image classification tasks in Figure 14, [^3] Image classification is the largest category in MLE-bench (8 tasks out of 22 in MLE-bench lite) we observe that AIDE relies on the EfficientNet architecture (and its variants) for almost 40% of the tasks. For 75% of its attempts, AIDE uses only 3 different architectures: EfficientNet, ResNet (He et al., 2016) and LightGBM. $\text{AIRA}_{\text{Greedy}}$ uses a wider range of architectures, with EfficientNet, ConvNeXt (Liu et al., 2022), and ViT (Dosovitskiy et al., 2020) making only 38% of the agents’ initial ideas.